OpenStudy (anonymous):

I WILL MEDAL BEST ANSWER What is the value of w? 5(9 − 3) = w • 9 – 5 • 3

OpenStudy (anonymous):

-3w=-5 ---divide both sides with -3 -3w/-3=-5/-3 w=5/3 (answer)

OpenStudy (madhu.mukherjee.946):

30 /12

OpenStudy (madhu.mukherjee.946):

or 15/6

OpenStudy (anonymous):

Okay, thank you guys very much, I wish I could could give a medal to all of you

OpenStudy (anonymous):

just click best response

OpenStudy (madhu.mukherjee.946):

5(6)=12w 30=12w w=30/12=15/6